OpenStudy (otonogold):

Determine whether y varies directly with x. If so, find the constant of variation k and write the equation. x y 4 6.4 7 11.2 10 16 13 20.8

OpenStudy (***[isuru]***):

If y is directly proportional to x , y should increase with a constant value when x is increasing with a constant value and vice and versa... hence the two values r keep increasing by constant values here x is getting increased by +3 each time and y keeps increasing by +4.8 as x increases therefor we can say y varies directly with x. hence x and y keep increasing from constant values we can say... y = kx and k = y/x = 6.4/4 therefor .... y =(6.4/4)x = (1.6)x multiply both sides by 10 10 y = 16x